# More is not always better? Enhancing Many-Shot In-Context Learning with Differentiated and Reweighting Objectives

### v2

@@ Abstract
Large language models (LLMs) excel at few-shot in-context learning (ICL) without requiring parameter updates. However, as the number of ICL demonstrations increases from a few to many, performance tends to plateau and eventually decline. We identify two primary causes for this trend: the suboptimal negative log-likelihood (NLL) optimization objective and the incremental data noise. To address these issues, we introduce DrICL, a novel optimization method that enhances model performance through Differentiated Learning and advantage-based Reweighting objectives. Globally, DrICL utilizes differentiated learning to optimize the NLL objective, ensuring that many-shot performance surpasses zero-shot levels. Locally, it dynamically adjusts the weighting of many-shot demonstrations by leveraging cumulative advantages inspired by reinforcement learning, thereby improving generalization. This approach allows the model to handle varying numbers of shots effectively, mitigating the impact of noisy data. Recognizing the lack of multi-task datasets with diverse many-shot distributions, we develop the Many-Shot ICL Benchmark (ICL-50)-a large-scale benchmark of 50 tasks that cover shot numbers from 1 to 350 within sequences of up to 8,000 tokens-for fine-tuning purposes. ICL-50 facilitates the evaluation of many-shot ICL strategies across seven prominent NLP tasks and 50 distinct datasets. Experimental results demonstrate that LLMs enhanced with DrICL achieve significant improvements in many-shot setups across various tasks, including both in-domain and out-of-domain scenarios. We release the code and benchmark dataset hoping to facilitate further research in many-shot ICL [^1] https://github.com/xiaoqzhwhu/DrICL.
@@
Xiaoqing Zhang1,2   Ang  Lv1   Yuhan  Liu1   Flood Sung2 Wei Liu3   Shuo Shang4   Xiuying Chen511footnotemark: 1     Rui  Yan1thanks:   Corresponding authors. 1Gaoling School of Artificial Intelligence, Renmin University of China  2MoonshotAI 3Xiaomi AI Lab 4University of Electronic Science and Technology of China 5Mohamed bin Zayed University of Artificial Intelligence {xiaoqingz, anglv, yuhan.liu, ruiyan}@ruc.edu.cn{floodsung}@moonshot.cn {liuwei40}@xiaomi.com{jedi.shang}@gmail.com{xy-chen}@pku.edu.cn
@@ 1 Introduction
In-context learning Brown et al. (2020) enables models to quickly adapt and address specific issues by utilizing contextual cues, improving adaptability and generalization. With the expansion of the context length in advanced LLMs, the ability to process text lengths up to 1 million tokens allows LLMs to accept increasingly more demonstrations. The ICL scenarios with hundreds or thousands of shots are called many-shot learning Agarwal et al. (2024). However, many-shot does not always result in better performance than few-shot. Some models exhibit a linear decrease in ICL capabilities with the increase in ultra-long text lengths Liu et al. (2024). As shown in Figure 1, we present the accuracy variations of Mistral-7B-Instruct-v0.2 on the CLSClusteringS2S dataset Li et al. (2022). As the number of ICL examples increases, models’ performance exhibits a trend of rising and then falling.
@@
To address the above factors, we propose DrICL, enhancing many-shot in-context learning with a refined fine-tuning objective. For the first factor, we propose differentiated learning to deal with the trade-off between many-shot and zero-shot scenarios from a global perspective. During differentiated learning, we ensure that the performance on many-shot demonstrations surpasses that on zero-shot demonstrations. This approach promotes model’s deeper understanding of contextual cues, encouraging the model to effectively leverage contextual information. For the second factor, inspired by reinforcement learning, we propose an advantage-based reweighting method to filter noise in many-shot demonstrations from a local perspective. In reinforcement learning, the “advantage function” is essential for assessing the value of actions beyond the average expected return, effectively directing the policy to choose actions that are predicted to yield the highest future rewards Baird (1994). We divide each sequence containing $K$ demonstrations into multiple reweighting windows, each with $W$ demonstrations. For each demonstration $x_{k}$ in the reweighting window $w$, we designate the preceding window $w-1$ as the sampling window. We calculate the cumulative advantage from the data within the sampling window $w-1$ and apply it dynamically as the reward for $x_{k}$. Then, we integrate the cumulative advantage into the NLL computation for $x_{k}$ and get an advantage-based training objective. We combine the global differentiated learning and local reweighting as the final refined training objective for maintaining ICL under many-shots.
@@
In summary, this paper identified two challenges in avoiding ICL’s decreasing performance under Many-shot scenarios: suboptimal training objective, and incremental data noise. We propose “differentiated learning” and “advantaged-based reweighting” to address these challenges, respectively. We further propose the largest ICL-50 to support our training, evaluation as well as further studies. We experiment DrICL on the ICL-50 with open-source LLMs, showing the stable performance both in-domain and out-of-domain under many-shots. All these points form the major contribution of this paper.
@@ 2 Related Work
Instruction Tuning of LLMs. Instruction tuning has become an effective technique for enhancing the capabilities and controllability of LLMs Zhang et al. (2023). In the domain of ICL, studies like MetaICL Min et al. (2022) and PEFT Bertsch et al. (2024) have demonstrated that fine-tuning LLMs with both small and large demonstration sizes, denoted as $k$, lead to improved ICL performance. Despite their studies being confined to a modest quantity of tuning data—capped at 10,000 entries—there is an evident necessity for deeper research into how ICL performs when scaled up with more extensive datasets. Consequently, we introduce ICL-50, a significantly larger dataset, designed to delve into the strategies for amplifying ICL’s potential.
@@ 3.2.1 Importance Sampling
Importance sampling adjusts the sampling weights to reduce bias and imbalance brought by noisy data. In this section, we leverage the loss of samples on the training dataset evaluated by the training model to calculate the importance weights of these samples. For each training sequence $S_{K}=\{x_{1}y_{1}x_{2}y_{2}\ldots x_{K}y_{K}\}$, we calculate the loss ${\mathcal{L}_{\text{many-shot}}}_{k}$ generated by the sequence $\{x_{1}y_{1}x_{2}y_{2}\ldots x_{k}\}$ at the current $k$-th position to represent the features of $S_{k}$. Our objective is to adjust the weighting of examples based on their significance, thereby ensuring that more emphasis is placed on the critical instances while reducing focus on the less important ones.
@@ 4.3.2 Performance Variance
We track the performance variance across all datasets with NFT(6.49E-03), IT(2.71E-02), MetaICL(2.38E-03), and DrICL(1.56E-03) as $k$ varied. On average, DrICL has the lowest variance, suggesting a more stable performance with $k$-shot of demonstration changes. For details of the performance variance of each dataset please refer to Table 10 in Appendix 8.
@@ 4.3.3 Data Noise Sensitivity
We compare DrICL with and without local reweighting by examining how loss variance trends for each $k$-shot demonstration during training. The reweighting window in DrICL reduces loss variance and effectively balances the impact of noisy data by appropriately weighting demonstrations. This reduction in sensitivity to data noise helps maintain stable performance as the number of demonstrations increases. For details of the noise variation please refer to Table 11 in Appendix 11.
@@ B.2 Hyperparameters Study
B.3 Performance Variance
Table 10 is the performance variance across the NFT, IT, MetaICL, and DrICL methods. In Table 10, we prove that our method demonstrates the smallest deviation, indicating greater stability in performance as $k$-shot changes.
Dataset NFT IT MetaICL DrICL
OpenbookQA 2.20E-03 3.90E-03 5.60E-03 8.00E-04
ARC 2.41E-02 2.10E-02 2.00E-03 1.40E-03
CLSClusteringS2S 2.40E-03 1.00E-02 3.00E-04 2.00E-04
ArxivClusteringS2S 1.00E-03 3.70E-03 8.00E-04 9.00E-04
TengkgnadClusteringS2S 7.00E-03 1.90E-03 5.00E-04 5.00E-04
TweetSentimentExtraction 3.30E-03 3.40E-03 4.90E-03 5.00E-04
GSM8K 8.50E-03 2.20E-03 1.00E-04 5.00E-04
XSUM 1.40E-03 2.20E-03 5.00E-05 5.00E-05
CNN 3.90E-03 2.30E-02 3.00E-03 3.70E-03
EcomRetrieval 4.10E-03 1.20E-01 7.00E-04 8.00E-04
VideoRetrieval 2.00E-02 1.10E-01 2.00E-03 0.00E+00
cMedQA 0.00E+00 2.40E-02 8.60E-03 9.40E-03
Average 6.49E-03 2.71E-02 2.38E-03 1.56E-03
Table 10: The performance variation of various datasets.
Methods Variance Trends Across Training Progress

### v3

@@ Abstract
Large language models (LLMs) excel at few-shot in-context learning (ICL) without requiring parameter updates. However, as ICL demonstrations increase from a few to many, performance tends to plateau and eventually decline. We identify two primary causes for this trend: the suboptimal negative log-likelihood (NLL) optimization objective and the incremental data noise. To address these issues, we introduce DrICL, a novel optimization method that enhances model performance through Differentiated and Reweighting objectives. Globally, DrICL utilizes differentiated learning to optimize the NLL objective, ensuring that many-shot performance surpasses zero-shot levels. Locally, it dynamically adjusts the weighting of many-shot demonstrations by leveraging cumulative advantages inspired by reinforcement learning, thereby mitigating the impact of noisy data. Recognizing the lack of multi-task datasets with diverse many-shot distributions, we develop the Many-Shot ICL Benchmark (ICL-50)-a large-scale benchmark of 50 tasks that cover shot numbers from 1 to 350 within sequences of up to 8,000 tokens-for both fine-tuning and evaluation purposes. Experimental results demonstrate that LLMs enhanced with DrICL achieve significant improvements in many-shot setups across various tasks, including both in-domain and out-of-domain scenarios. We release the code and dataset hoping to facilitate further research in many-shot ICL [^1] https://github.com/xiaoqzhwhu/DrICL.
@@
Xiaoqing Zhang1,2thanks: This work was done during the internship at Moonshot AI.   Ang  Lv1   Yuhan  Liu1   Flood Sung2 Wei Liu3   Jian Luan3   Shuo Shang4   Xiuying Chen522footnotemark: 2     Rui  Yan1,6,7thanks:   Corresponding authors. 1Gaoling School of Artificial Intelligence, Renmin University of China  2MoonshotAI  3Xiaomi AI Lab 4University of Electronic Science and Technology of China  5Mohamed bin Zayed University of Artificial Intelligence 6School of Artifcial Intelligence, Wuhan University 7Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MoE {xiaoqingz, anglv, yuhan.liu, ruiyan}@ruc.edu.cnfloodsung@moonshot.cn {liuwei40,luanjian}@xiaomi.comjedi.shang@gmail.comxy-chen@pku.edu.cn
@@ 1 Introduction
In-context learning Brown et al. (2020) enables models to quickly adapt and address specific issues by utilizing contextual cues, improving adaptability and generalization. With the expansion of the context length in advanced LLMs, the ability to process text lengths up to 1 million tokens allows LLMs to accept increasingly more demonstrations. The ICL scenarios with hundreds or thousands of shots are called many-shot learning Agarwal et al. (2024). However, many-shot does not always result in better performance than few-shot. Some models exhibit a linear decrease in ICL capabilities with the increase in ultra-long text lengths Liu et al. (2024a). As shown in Figure 1, we present the accuracy variations of Mistral-7B-Instruct-v0.2 on the CLSClusteringS2S dataset Li et al. (2022). As the number of ICL examples increases, models’ performance exhibits a trend of rising and then falling.
@@
To address the above factors, we propose DrICL, enhancing many-shot in-context learning with a refined fine-tuning objective. For the first factor, we propose differentiated learning to deal with the trade-off between many-shot and zero-shot scenarios from a global perspective. During differentiated learning, we ensure that the performance on many-shot demonstrations surpasses that on zero-shot demonstrations. This approach promotes the model’s deeper understanding of contextual cues, encouraging it to leverage contextual information effectively. For the second factor, we find the noise in the demonstrations is reflected in the sharp increase in loss observed for certain samples, which disrupts the training process. Inspired by reinforcement learning, we propose an advantage-based reweighting method to reduce focus on noise in many-shot demonstrations from a local perspective. In reinforcement learning, the “advantage function” is essential for assessing the value of actions beyond the average expected return, effectively directing the policy to choose actions that are predicted to yield the highest future rewards Baird (1994). Similarly, just as the advantage function helps identify actions that yield higher returns than the average, we extend this concept by using cumulative advantage to adjust the weights of demonstrations. This adjustment ensures that demonstrations with extreme loss fluctuations do not disproportionately influence the model or disrupt stable learning. The cumulative advantage for each demonstration is calculated based on the loss of the current demonstration and the losses of sampled demonstrations from preceding ones. We introduce a sampling window to ensure a balanced consideration of previous demonstrations. Each sequence is divided into multiple reweighting windows, and for each demonstration in a reweighting window $w$, the preceding window $w-1$ serves as the sampling window. Finally, we incorporate the cumulative advantage into the negative log-likelihood (NLL) computation, resulting in an advantage-based training objective.
@@
In summary, this paper identified two challenges in avoiding ICL’s decreasing performance under Many-shot scenarios: suboptimal training objective, and incremental data noise. We propose “differentiated learning” and “advantaged-based reweighting” to address these challenges, respectively. We further propose the largest ICL-50 to support our training, evaluation as well as further studies. We experiment DrICL on the ICL-50 with open-source LLMs, showing stable performance both in-domain and out-of-domain under many-shots. All these points form the major contribution of this paper.
@@ 2 Related Work
Instruction Tuning of LLMs. Instruction tuning has become an effective technique for enhancing the capabilities and controllability of LLMs Zhang et al. (2023). In the domain of ICL, studies like MetaICL Min et al. (2022) , IAD Liu et al. (2024b), and PEFT Bertsch et al. (2024) have demonstrated that fine-tuning LLMs with both small and large demonstration sizes, denoted as $k$, lead to improved ICL performance. Despite their studies being confined to a modest quantity of tuning data—capped at 10,000 entries—there is an evident necessity for deeper research into how ICL performs when scaled up with more extensive datasets. Consequently, we introduce ICL-50, a significantly larger dataset, designed to delve into the strategies for amplifying ICL’s potential.
@@ 3.2.1 Importance Sampling
Importance sampling adjusts weights to reduce bias and imbalance from noisy data. Here, we use the training model’s loss on samples to calculate their importance weights. For each training sequence $S_{K}=\{x_{1}y_{1}x_{2}y_{2}\ldots x_{K}y_{K}\}$, we calculate the loss ${\mathcal{L}_{\text{many-shot}}}_{k}$ generated by the sequence $\{x_{1}y_{1}x_{2}y_{2}\ldots x_{k}\}$ at the current $k$-th position to represent the features of $S_{k}$. Our goal is to weight examples by their significance, emphasizing critical instances and reducing focus on less important ones.
@@ 4.3.1 Performance Tradeoff
Dataset NFT IT MetaICL DrICL
OpenbookQA 2.20E-03 3.90E-03 5.60E-03 8.00E-04
ARC 2.41E-02 2.10E-02 2.00E-03 1.40E-03
CLSClusteringS2S 2.40E-03 1.00E-02 3.00E-04 2.00E-04
ArxivClusteringS2S 1.00E-03 3.70E-03 8.00E-04 9.00E-04
TengkgnadClusteringS2S 7.00E-03 1.90E-03 5.00E-04 5.00E-04
TweetSentimentExtraction 3.30E-03 3.40E-03 4.90E-03 5.00E-04
GSM8K 8.50E-03 2.20E-03 1.00E-04 5.00E-04
XSUM 1.40E-03 2.20E-03 5.00E-05 5.00E-05
CNN 3.90E-03 2.30E-02 3.00E-03 3.70E-03
EcomRetrieval 4.10E-03 1.20E-01 7.00E-04 8.00E-04
VideoRetrieval 2.00E-02 1.10E-01 2.00E-03 0.00E+00
cMedQA 0.00E+00 2.40E-02 8.60E-03 9.40E-03
Average 6.49E-03 2.71E-02 2.38E-03 1.56E-03
Table 7: The performance variation of various datasets.
@@ 4.3.2 Performance Variance
Table 7 is the performance variance across the NFT, IT, MetaICL, and DrICL methods. We track the performance variance across all datasets with NFT(6.49E-03), IT(2.71E-02), MetaICL(2.38E-03), and DrICL(1.56E-03) as $k$ varied. We prove that our method demonstrates the smallest deviation, indicating greater stability in performance as $k$-shot changes.
@@ 4.3.3 Data Noise Sensitivity
We compare DrICL with and without local reweighting by examining how loss variance trends for each $k$-shot demonstration during training. The reweighting window in DrICL reduces loss variance and effectively balances the impact of noisy data by appropriately weighting demonstrations. This reduction in sensitivity to data noise helps maintain stable performance as the number of demonstrations increases. For details of the noise variation please refer to Table 11 in Appendix B.2.
@@ 5 Conclusions
Acknowledgments
This work is also supported by the Public Computing Cloud, Renmin University of China and by fund for building worldclass universities (disciplines) of Renmin University of China.
Limitations
In this work, we balance the samples in the training set, but we have not fully analyzed the algorithm’s robustness across datasets of varying sizes. As a result, DrICL’s performance may vary when applied to datasets of different scales, which we plan to explore in future work. Regarding window size design, we used a uniform size for all samples. However, tasks with varying sample lengths may result in oversampling for short-text tasks and undersampling for long-text tasks. To address this, we plan to implement dynamic window sizes based on sample lengths to ensure balanced representation for both short and long samples.
@@ B.2 Hyperparameters Study
B.3 Data Noise Sensitivity
Methods Variance Across Training Progress